Solve the equation: x - 5 = the square root of 4x - 15
Answer by mananth(16946)

x - 5 = sqrt(4x - 15)
square both sides
(x-5)^2= 4x-15
x^2-10x+25=4x-15
x^2-10x-4x=-25-15
x^2-14x+40=0
x^2-10x-4x+40=0
x(x-10)-4(x-10)=0
(x-10)(x-4)=0
x= 10 OR x =4